# Provable Privacy with Non-Private Pre-Processing

### v1

@@ A.2.2 Gradient-based methods
In this section, we start with providing the formal definitions of gradient-based methods mentioned in Section 2.2 and evaluated in Table 1, including DP-GD, DP-SGD with subsampling, and DP-SGD with iteration. Then, we provide the detailed theorems and proofs for their Smooth RDP parameters, as stated in Table 1, in Theorem 12, 13 and 14.
@@
Given a $L$-Lipschitz loss function $\ell$, DP-GD [Bassily et al., 2014, Song et al., 2021], denoted by ${\mathcal{A}}_{\mathrm{DP-GD}}$ starts from some random initialization $w_{0}$ in the parameter space ${\mathcal{H}}$ and conducts projected gradient descent for $T$ iterations as $w_{t+1}=\Pi_{{\mathcal{H}}}(w_{t}-\eta\tilde{g}_{t})$ with the noisy gradient on the whole dataset $\tilde{g}_{t}$ defined as $\tilde{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{w}\ell(w_{t-1},x)+{\mathcal{N}}(0,\sigma^{2})$, and outputs the average parameter over the $T$ iterations $\frac{1}{T}\sum_{t=1}^{T}w_{t}$.
@@
Another commonly used method is DP-SGD with subsampling [Abadi et al., 2016, Bassily et al., 2014], denoted as ${\mathcal{A}}_{\mathrm{SGD-samp}}$. In each gradient descent step, DP-SGD with subsampling first draws a uniform subsample of size $B$ from the dataset without replacement. The gradient update is then performed by adding noise to the average gradient derived from this subsample. In contrast, DP-GD computes the average gradient of the entire dataset for its updates.
@@
Differentially Private Fixed Gradient Descent (DP-FGD), denoted by ${\mathcal{A}}_{\mathrm{FGD}}^{(T)}$, is a variant of DP-SGD. It processes the data points in a fixed order — the gradient at the $t^{\text{th}}$ step is calculates using the $t^{\text{th}}$ point in the dataset— and outputs the parameter obtained after the $T^{\text{th}}$ iteration. DP-SGD with iteration [Feldman et al., 2018], denoted as ${\mathcal{A}}_{\mathrm{SGD-iter}}$, uses DP-FGD as a base procedure. It takes an extra parameter $\kappa_{\tau}$ and first uniformly samples an integer $T$ from $\left[{n-\kappa_{\tau}+1}\right]=\left\{{1,\ldots,n-\kappa_{\tau}+1}\right\}$ and releases the output from ${\mathcal{A}}_{\mathrm{FGD}}^{(T)}$, i.e., it releases the result from the $T^{\text{th}}$ iteration. While DP-SGD with iteration cannot take advantage of privacy amplification by subsampling for its privacy analysis, it relies on privacy amplification with iteration [Feldman et al., 2018] to achieve a comparable privacy guarantee to that of ${\mathcal{A}}_{\mathrm{SGD-samp}}$.
@@
Lemma E (Post-processing of RDP [Mironov, 2017]).
@@
Lemma F (Composition of RDP, Proposition 1 in Mironov [2017]).
@@
Lemma G (Lemma 3 in Abadi et al. [2016]).
@@
We first note that ${\mathcal{A}}_{\mathrm{SGD-iter}}$ is $(\alpha,\frac{\alpha\varepsilon^{2}}{2})$-RDP following Theorem 26 in Feldman et al. [2018] (see H below for completeness).
Lemma H (Privacy guarantee of SGD by iteration, Theorem 26 in Feldman et al. [2018]).
@@
The functions $\psi_{i}$ and $\psi_{i}^{\prime}$ are contractive functions for $\eta\leq 2/\beta$ [Nesterov, 2004]. It follows by the definition of DPFGD that $\theta_{T}={\mathcal{A}}_{\mathrm{FGD}}^{T}(S)$ and $\theta_{T}^{\prime}={\mathcal{A}}_{\mathrm{FGD}}^{T}(S^{\prime})$ are the $T^{\text{th}}$ outputs of the CNIs $(\theta_{0},\left\{{\psi_{t}}\right\},{\mathcal{N}}(0,\left({\eta\sigma}\right)^{2}))$ and $(\theta_{0},\left\{{\psi_{t}^{\prime}}\right\},{\mathcal{N}}(0,\left({\eta\sigma}\right)^{2}))$ respectively.
@@
Lemma I (Theorem 22 in Feldman et al. [2018] with fixed noise distribution).
@@ Appendix B Proof of Meta-Theorem (Theorem 2)
Lemma J (Triangle inequality of Rényi divergence Mironov [2017]).
@@
Using J, for any $p>1$, we have
$D_{\alpha}({\mathcal{A}}(\pi_{1}(S_{1}))||{\mathcal{A}}(\pi_{2}(S_{2})))$ $\leq\frac{\alpha-\frac{1}{p}}{\alpha-1}D_{p\alpha}({\mathcal{A}}(\pi_{1}(S_{1}))|{\mathcal{A}}(\tilde{S}))+D_{\frac{p\alpha-1}{p-1}}({\mathcal{A}}(\tilde{S})||{\mathcal{A}}(\pi_{2}(S_{2})))$ (30)
$\overset{(a)}{\leq}\frac{\alpha p-1}{p(\alpha-1)}\tilde{\varepsilon}(\alpha p,\Delta_{\infty}\Delta_{2})+\varepsilon\left({\frac{p\alpha-1}{p-1}}\right)$
@@
Similarly, by constructing a dataset $\tilde{S}$ consisting of $\pi_{1}(S)$ and $\pi_{2}(z_{2})$, we upper bound $D_{\alpha}\left({{\mathcal{A}}(\pi_{2}(S_{2}))}||{{\mathcal{A}}(\tilde{S})}\right)$ and $D_{\alpha}\left({{\mathcal{A}}(\tilde{S})}||{{\mathcal{A}}(\pi_{1}(S_{1}))}\right)$ with the SRDP and RDP property in a similar manner as Equations 28 and 29. Applying J, we can show that for any $q>1$,
$D_{\alpha}({\mathcal{A}}(\pi_{2}(S_{2}))||{\mathcal{A}}(\pi_{1}(S_{1})))\leq\frac{\alpha q-1}{q(\alpha-1)}\varepsilon(\alpha q)+\tilde{\varepsilon}\left({\frac{q\alpha-1}{q-1},\Delta_{\infty}\Delta_{2}}\right)$ (31)
@@ C.2 Sensitivity analysis of model-based imputation
Many imputation methods, such as median and linear regression, do not have bounded global sensitivity even when the instance space is bounded [Alabi et al., 2022]. However, the local sensitivity of these methods are usually bounded on well-behaved datasets. Given a collection of well-behaved datasets ${\mathcal{L}}$, the $L_{2}$ sensitivity over the collection is the upper bound on the local sensitivity of all datasets $S\in{\mathcal{L}}$. We exploit this property to find the $L_{2}$ sensitivity of the aforementioned imputation methods. We introduce additional notation and provide a summary of the $L_{2}$ sensitivity for different methods in Table 3.
@@
[…]
Notation
Meaning
Imputation Model $f$
$L_{2}$-Sensitivity $\Delta_{2}({\mathcal{L}},\pi_{f})$
$p$
Maximum number of missing points in any $S\in\mathcal{L}$
Mean
$\frac{2}{n-p}$
$x_{(m)}^{\min}$
Minimum $m^{th}$ ordered statistics of any feature in any $S\in\mathcal{L}$
Median
$x_{\left({(n+1)/2}\right)}^{\max}-x_{\left({n/2}\right)}^{\min}$
$x_{(m)}^{\max}$
Maximum $m^{th}$ ordered statistics of any feature in any $S\in\mathcal{L}$
$m$-Trimmed Mean
$\nicefrac{{x_{(m)}^{\max}-x_{(n-m)}^{\min}}}{{n-2m-p}}$
$\lambda_{\max}$
Maximum eigenvalue of $X^{T}X$, for any $S=(X,Y)\in\mathcal{L}$
Linear regression
$\frac{\lambda_{\max}^{2}}{(\lambda_{\max}+1)\lambda_{\min}^{2}}+\frac{1}{\lambda_{\min}}$
$\lambda_{\min}$
Minimum eigenvalue of $X^{T}X$, for any $S=(X,Y)\in\mathcal{L}$
@@
where step (a) follows from Sherman-Morrison-Formula and that the eigenvalue of a principal submatrix is always larger than the smallest eigenvalue of the original matrix following Theorem 4.3.15 in Horn and Johnson [1985] and then taking supremum over all $X\in{\mathcal{L}}$. Similarly, step (b) follows from the fact that the eigenvalue of a principal submatrix is always smaller than the largest eigenvalue of the original matrix following Theorem 4.3.15 in Horn and Johnson [1985] and then taking supremum over all $X\in{\mathcal{L}}$. ∎
@@ C.3 Sensitivity analysis of PCA
Lemma K (Davis-Kahan Theorem [Yu et al., 2014]).
@@
=$\overset{(a)}{=}\underbrace{\frac{1}{n-1}\sum_{i=1}^{n-1}(x_{i}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))(x_{i}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))^{\top}}_{\text{part I}}+\underbrace{\frac{1}{n-1}(x_{n}^{\prime}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))(x_{n}^{\prime}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))^{\top}}_{\text{part II}}$
@@
Lemma L (Simplified version of Theorem 3 in Zwald and Blanchard [2005]).
@@
In this section, we provide the proofs of the overall privacy guarantees for specific pre-processed DP pipelines, as stated in Table 2 in Section 4.2. For clarity, we restate the full version of Theorem 6 and specify the privacy guarantee for each category of privacy mechanisms (each row in Table 2) in Theorem 16. Then, we present the proof for each category separately.
@@
    Gaussian mechanism with output function $f$ satisfies $\left({\alpha,{1.05\alpha\varepsilon^{2}}(1+p^{2})}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}(1+\eta^{2}p^{2})}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}}{(n-p)^{2}}}\right)}\right)$-RDP and $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({1+\frac{12.2^{2}}{\delta_{\min}^{2}}}\right)}\right)$-RDP when coupled with deduplication, quantization, mean imputation and PCA for rank reduction respectively. DP-GD with loss function $\ell$ satisfies $\left({\alpha,{1.05\alpha\varepsilon^{2}}(4+p^{2})}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}(4+\eta^{2}p^{2})}\right)$-RDP, $\left({\alpha,{4.2\alpha\varepsilon^{2}}\left({1+\frac{p^{2}}{(n-p)^{2}}}\right)}\right)$-RDP and $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({4+\frac{12.2^{2}}{\delta_{\min}^{2}}}\right)}\right)$-RDP when coupled with deduplication, quantization, mean imputation and PCA for rank reduction respectively.
@@
    Exponential mechanism with score function $Q$ and Laplace mechanism with output function $f$ satisfy $(\alpha,\varepsilon(1+p))$-RDP, $(\alpha,\varepsilon(1+\eta p))$-RDP, $\left({\alpha,\varepsilon\left({1+\frac{2p}{n-p}}\right)}\right)$-RDP and $\left({\alpha,\varepsilon\left({1+\frac{12.2}{\delta{\min}}}\right)}\right)$-RDP when coupled with deduplication, quantization, mean imputation and PCA for rank reduction respectively.
@@
    DP-SGD with subsampling with loss function $\ell$ satisfies $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({2\alpha+\frac{12.2^{2}}{\delta_{\min}^{2}}}\right)}\right)$-RDP when coupled with PCA for rank reduction .
@@
We apply Theorem 2 with $p=q=2$,
@@
By substituting the expression of RDP and SRDP parameter from Table 1 for Gaussian mechanism, the $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for deduplication, quantization, imputation and PCA from Proposition 3Corollary 4 and Proposition 5, and the Lipschitz parameter and global sensitivity of the output function $f$ into Equation 43, we complete the proof of the overall privacy guarantees for Gaussian mechanism.
Similarly, for DP-GD, we substitute the expression of RDP and SRDP parameter from Table 1 for DP-GD, the $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for deduplication, quantization, imputation and PCA from Proposition 3Corollary 4 and Proposition 5, and the Lipschitz and smoothness parameter of the loss function $\ell$ into Equation 43. This completes the proof of the overall privacy guarantees for DP-GD.
@@
We apply Theorem 2 with $p=q=1$,
@@
Similarly, for exponential mechanism, we substitute the expression of RDP and SRDP parameter from Table 1 for exponential mechanism, the $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for deduplication, quantization, imputation and PCA from Proposition 3Corollary 4 and Proposition 5, and the Lipschitzness and the global sensitivity of the score function $Q$, into Equation 43. This completes the proof of the overall privacy guarantees for exponential mechanism.
@@
Following a similar argument as the proof for DP-GD, we set $p=q=2$ in Theorem 2 and obtain Equation 43. Then, we substitute in Equation 43 the following parameters: a)RDP and SRDP parameters for DP-SGD with subsampling from Table 1, b)$\gamma=1$, c) Lipschitzness and smoothness parameter of the loss function $\ell$, and d) $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for PCA from Proposition 5. This completes the proof. ∎
For DP-SGD with iteration, the pre-processing mechanisms achieve a tighter bound while satisfying $\Delta_{\infty}\ll n$ by Definition 8. Therefore, this method will not provide tighter bound for PCA. However, it improves the privacy analysis for imputation, deduplication and quantization, as proved below.
@@
Following a similar argument as the proof for DP-GD, we set $p=q=2$ in Theorem 2. We substitute in Equation 43 the following parameters: a)RDP and SRDP parameters for DP-SGD with iteration from Table 1, b)$\kappa_{\tau}=\Delta_{\infty}=p$, c) Lipschitzness and smoothness parameter of the loss function $\ell$, and d) $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for quantization, deduplication and imputation from from Proposition 3 and Corollary 4. This completes the proof. ∎

### v4

@@ A.2.2 Gradient-based methods
In this section, we start with providing the formal definitions of gradient-based methods mentioned in Section 2.2 and evaluated in Table 1, including DP-GD, DP-SGD with subsampling, and DP-SGD with iteration. Then, we provide the detailed theorems and proofs for their Smooth RDP parameters, as stated in Table 1, in Theorem 13, 14 and 15.
@@
Given a $L$-Lipschitz loss function $\ell$, DP-GD (Bassily et al., 2014; Song et al., 2021), denoted by ${\mathcal{A}}_{\mathrm{DP-GD}}$ starts from some random initialization $w_{0}$ in the parameter space ${\mathcal{H}}$ and conducts projected gradient descent for $T$ iterations as $w_{t+1}=\Pi_{{\mathcal{H}}}(w_{t}-\eta\tilde{g}_{t})$ with the noisy gradient on the whole dataset $\tilde{g}_{t}$ defined as $\tilde{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\nabla_{w}\ell(w_{t-1},x)+{\mathcal{N}}(0,\sigma^{2})$, and outputs the average parameter over the $T$ iterations $\frac{1}{T}\sum_{t=1}^{T}w_{t}$.
@@
Another commonly used method is DP-SGD with subsampling (Abadi et al., 2016; Bassily et al., 2014), denoted as ${\mathcal{A}}_{\mathrm{SGD-samp}}$. In each gradient descent step, DP-SGD with subsampling first draws a uniform subsample of size $B$ from the dataset without replacement. The gradient update is then performed by adding noise to the average gradient derived from this subsample. In contrast, DP-GD computes the average gradient of the entire dataset for its updates.
@@
Differentially Private Fixed Gradient Descent (DP-FGD), denoted by ${\mathcal{A}}_{\mathrm{FGD}}^{(T)}$, is a variant of DP-SGD. It processes the data points in a fixed order — the gradient at the $t^{\text{th}}$ step is calculates using the $t^{\text{th}}$ point in the dataset— and outputs the parameter obtained after the $T^{\text{th}}$ iteration. DP-SGD with iteration (Feldman et al., 2018), denoted as ${\mathcal{A}}_{\mathrm{SGD-iter}}$, uses DP-FGD as a base procedure. It takes an extra parameter $\kappa_{\tau}$ and first uniformly samples an integer $T$ from $\left[{n-\kappa_{\tau}+1}\right]=\left\{{1,\ldots,n-\kappa_{\tau}+1}\right\}$ and releases the output from ${\mathcal{A}}_{\mathrm{FGD}}^{(T)}$, i.e., it releases the result from the $T^{\text{th}}$ iteration. While DP-SGD with iteration cannot take advantage of privacy amplification by subsampling for its privacy analysis, it relies on privacy amplification with iteration (Feldman et al., 2018) to achieve a comparable privacy guarantee to that of ${\mathcal{A}}_{\mathrm{SGD-samp}}$.
@@
Lemma E (Post-processing of RDP (Mironov, 2017)).
@@
Lemma F (Composition of RDP, Proposition 1 in Mironov (2017)).
@@
Lemma G (Lemma 3 in Abadi et al. (2016)).
@@
We first note that ${\mathcal{A}}_{\mathrm{SGD-iter}}$ is $(\alpha,\frac{\alpha\varepsilon^{2}}{2})$-RDP following Theorem 26 in Feldman et al. (2018) (see H below for completeness).
Lemma H (Privacy guarantee of SGD by iteration, Theorem 26 in Feldman et al. (2018)).
@@
The functions $\psi_{i}$ and $\psi_{i}^{\prime}$ are contractive functions for $\eta\leq 2/\beta$ (Nesterov, 2004). It follows by the definition of DPFGD that $\theta_{T}={\mathcal{A}}_{\mathrm{FGD}}^{T}(S)$ and $\theta_{T}^{\prime}={\mathcal{A}}_{\mathrm{FGD}}^{T}(S^{\prime})$ are the $T^{\text{th}}$ outputs of the CNIs $(\theta_{0},\left\{{\psi_{t}}\right\},{\mathcal{N}}(0,\left({\eta\sigma}\right)^{2}))$ and $(\theta_{0},\left\{{\psi_{t}^{\prime}}\right\},{\mathcal{N}}(0,\left({\eta\sigma}\right)^{2}))$ respectively.
@@
Lemma I (Theorem 22 in Feldman et al. (2018) with fixed noise distribution).
@@ Appendix B Proof of Meta-Theorem (Theorem 2)
Lemma J (Triangle inequality of Rényi divergence Mironov (2017)).
@@
Using J, for any $c_{1}>1$, we have
$D_{\alpha}({\mathcal{A}}(\pi_{1}(S_{1}))||{\mathcal{A}}(\pi_{2}(S_{2})))$ $\leq\frac{\alpha-\frac{1}{c_{1}}}{\alpha-1}D_{c_{1}\alpha}({\mathcal{A}}(\pi_{1}(S_{1}))|{\mathcal{A}}(\tilde{S}))+D_{\frac{c_{1}\alpha-1}{c_{1}-1}}({\mathcal{A}}(\tilde{S})||{\mathcal{A}}(\pi_{2}(S_{2})))$ (30)
$\overset{(a)}{\leq}\frac{\alpha c_{1}-1}{c_{1}(\alpha-1)}\tilde{\varepsilon}(\alpha c_{1},\Delta_{\infty}\Delta_{2})+\varepsilon\left({\frac{c_{1}\alpha-1}{c_{1}-1}}\right)$
@@
Similarly, by constructing a dataset $\tilde{S}$ consisting of $\pi_{1}(S)$ and $\pi_{2}(z_{2})$, we upper bound $D_{\alpha}\left({{\mathcal{A}}(\pi_{2}(S_{2}))}||{{\mathcal{A}}(\tilde{S})}\right)$ and $D_{\alpha}\left({{\mathcal{A}}(\tilde{S})}||{{\mathcal{A}}(\pi_{1}(S_{1}))}\right)$ with the SRDP and RDP property in a similar manner as Equations 28 and 29. Applying J, we can show that for any $c_{2}>1$,
$D_{\alpha}({\mathcal{A}}(\pi_{2}(S_{2}))||{\mathcal{A}}(\pi_{1}(S_{1})))\leq\frac{\alpha c_{2}-1}{c_{2}(\alpha-1)}\varepsilon(\alpha c_{2})+\tilde{\varepsilon}\left({\frac{c_{2}\alpha-1}{c_{2}-1},\Delta_{\infty}\Delta_{2}}\right)$ (31)
@@ C.2 Sensitivity analysis of model-based imputation
Many imputation methods, such as median and linear regression, do not have bounded global sensitivity even when the instance space is bounded (Alabi et al., 2022). However, the local sensitivity of these methods are usually bounded on well-behaved datasets. Given a collection of well-behaved datasets ${\mathcal{L}}$, the $L_{2}$ sensitivity over the collection is the upper bound on the local sensitivity of all datasets $S\in{\mathcal{L}}$. We exploit this property to find the $L_{2}$ sensitivity of the aforementioned imputation methods. We introduce additional notation and provide a summary of the $L_{2}$ sensitivity for different methods in Table 3.
@@
where $x_{j}$ denotes the $j^{\text{th}}$ feature of $x$. Specifically, for mean imputation, median imputation, trimmed mean imputation, and regression, we present their sensitivities and corresponding assumptions in Table 3.
Notation Meaning Imputation Model $f$ $L_{2}$-Sensitivity $\Delta_{2}({\mathcal{L}},\pi_{f})$
$p$ Maximum number of missing points in any $S\in\mathcal{L}$ Mean $\frac{2}{n-p}$
$x_{(m)}^{\min}$ Minimum $m^{th}$ ordered statistics of any feature in any $S\in\mathcal{L}$ Median $x_{\left({(n+1)/2}\right)}^{\max}-x_{\left({n/2}\right)}^{\min}$
$x_{(m)}^{\max}$ Maximum $m^{th}$ ordered statistics of any feature in any $S\in\mathcal{L}$ $m$-Trimmed Mean $\nicefrac{{x_{(m)}^{\max}-x_{(n-m)}^{\min}}}{{n-2m-p}}$
$\lambda_{\max}$ Maximum eigenvalue of $X^{T}X$, for any $S=(X,Y)\in\mathcal{L}$ Linear regression $\frac{\lambda_{\max}^{2}}{(\lambda_{\max}+1)\lambda_{\min}^{2}}+\frac{1}{\lambda_{\min}}$
$\lambda_{\min}$ Minimum eigenvalue of $X^{T}X$, for any $S=(X,Y)\in\mathcal{L}$
@@
where step (a) follows from Sherman-Morrison-Formula and that the eigenvalue of a principal submatrix is always larger than the smallest eigenvalue of the original matrix following Theorem 4.3.15 in Horn & Johnson (1985) and then taking supremum over all $X\in{\mathcal{L}}$. Similarly, step (b) follows from the fact that the eigenvalue of a principal submatrix is always smaller than the largest eigenvalue of the original matrix following Theorem 4.3.15 in Horn & Johnson (1985) and then taking supremum over all $X\in{\mathcal{L}}$. ∎
@@ C.3 Sensitivity analysis of PCA
Lemma K (Davis-Kahan Theorem (Yu et al., 2014)).
@@
=$\overset{(a)}{=}$ $\underbrace{\frac{1}{n-1}\sum_{i=1}^{n-1}(x_{i}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))(x_{i}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))^{\top}}_{\text{part I}}$
$+\underbrace{\frac{1}{n-1}(x_{n}^{\prime}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))(x_{n}^{\prime}-\mu-\frac{1}{n}\left({x_{n}^{\prime}-x_{n}}\right))^{\top}}_{\text{part II}}$
@@
Lemma L (Simplified version of Theorem 3 in Zwald & Blanchard (2005)).
@@
C.4 Sensitivity Analysis of Scaling
See 6
Proof.
The $L_{\infty}$ sensitivity for both scaling methods is trivially upper bounded by the size of the datasets in ${\mathcal{L}}$. In the following, we prove the $L_{2}$ sensitivity for standard scaling and min max scaling respectively.
Proof of $L_{2}$ sensitivity for standard scaling
For any two neighboring datasets $S,S^{\prime}$, let $\mu,\mu^{\prime}$ denote the mean of $S,S^{\prime}$ respectively and $\sigma,\sigma^{\prime}$ denote the standard deviation of $S,S^{\prime}$ respectively. Then, the $L_{2}$ sensitivity of standard scaling is
$\Delta_{2}$ $=\max_{S,S^{\prime},x}\left\lVert\frac{x-\mu}{\sigma}-\frac{x-\mu^{\prime}}{\sigma^{\prime}}\right\rVert_{2}$ (43)
$=\max_{S,S^{\prime},x}\left\lVert\frac{\sigma^{\prime}x-\sigma^{\prime}\mu-\sigma x+\sigma\mu^{\prime}}{\sigma\sigma^{\prime}}\right\rVert$
$=\max_{S,S^{\prime},x}\left\lVert\frac{\left({\sigma^{\prime}-\sigma}\right)(x-\mu)}{\sigma\sigma^{\prime}}\right\rVert_{2}+\max_{S,S^{\prime}}\left\lVert\frac{\mu-\mu^{\prime}}{\sigma^{\prime}}\right\rVert$
$\overset{(a)}{\leq}\frac{2\max_{S,S^{\prime}}\left\lVert\sigma^{\prime}-\sigma\right\rVert}{\sigma_{\min}^{2}}+\frac{\max_{S,S^{\prime}}\left\lVert\mu-\mu^{\prime}\right\rVert}{\sigma_{\min}}$
where step (a) follows from $\left\lVert x-\mu\right\rVert_{2}\leq 2$ for any $x$ and $S$ and the definition of $\sigma_{\min}$.
From Liu (2016), the global sensitivity of sample mean and variance are $\frac{2}{n}$. We then show that for a dataset collection ${\mathcal{L}}$, the global sensitivity conditional on ${\mathcal{L}}$ is upper bounded by $\nicefrac{{1}}{{\sigma_{\min}n}}$.
For any $S,S^{\prime}$ with sample variances $\sigma^{2},\left({\sigma^{\prime}}\right)^{2}$, we have
$\left|{\sigma-\sigma^{\prime}}\right|\leq\frac{\left|{\sigma^{2}-\left({\sigma^{\prime}}\right)^{2}}\right|}{2\sigma_{\min}}\leq\frac{1}{\sigma_{\min}n},$ (44)
where the first inequality follows by rearranging $\left|{\sigma^{2}-\left({\sigma^{\prime}}\right)^{2}}\right|=\left|{\left({\sigma-\sigma^{\prime}}\right)\left({\sigma+\sigma^{\prime}}\right)}\right|\geq 2\sigma_{\min}\left|{\sigma-\sigma^{\prime}}\right|$, and the second inequality by substituting the global sensitivity of sample covariance from Liu (2016).
Substituting the sensitivity of sample mean $2/n$ and sample standard deviation $\frac{1}{\sigma_{\min}n}$ into Equation 43, we conclude the proof. ∎
@@ C.4 Sensitivity Analysis of Scaling
In this section, we provide the proofs of the overall privacy guarantees for specific pre-processed DP pipelines, as stated in Table 2 in Section 4.2. For clarity, we restate the full version of Theorem 7 and specify the privacy guarantee for each category of privacy mechanisms (each row in Table 2) in Theorem 17. Then, we present the proof for each category separately. As the proof of DP-GD is the same as that of Gaussian mechanism, we omit the proof for DP-GD for simplicity.
@@
    Gaussian mechanism with output function $f$ satisfy $\left({\alpha,{1.05\alpha\varepsilon^{2}}(1+p^{2})}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}(1+\eta^{2}p^{2})}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({1+\frac{4p^{2}}{(n-p)^{2}}}\right)}\right)$-RDP, $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({1+\frac{12.2^{2}}{\delta_{\min}^{2}}}\right)}\right)$-RDP and $1.05\alpha\varepsilon^{2}\left({1+\frac{4}{\sigma_{\min}^{3}}}\right)$-RDP when coupled with deduplication, quantization, mean imputation, PCA for rank reduction and standard scaling respectively.
@@
    Exponential mechanism with score function $Q$ and Laplace mechanism with output function $f$ satisfy $(\alpha,\varepsilon(1+p))$-RDP, $(\alpha,\varepsilon(1+\eta p))$-RDP, $\left({\alpha,\varepsilon\left({1+\frac{2p}{n-p}}\right)}\right)$-RDP, $\left({\alpha,\varepsilon\left({1+\frac{12.2}{\delta{\min}}}\right)}\right)$-RDP and $4.2\alpha\varepsilon^{2}\left({1+\frac{1}{\sigma_{\min}^{3}}}\right)$-RDP when coupled with deduplication, quantization, mean imputation, PCA and standard scaling for rank reduction respectively.
@@
    DP-SGD with subsampling with loss function $\ell$ satisfies $\left({\alpha,{1.05\alpha\varepsilon^{2}}\left({2\alpha+\frac{12.2^{2}}{\delta_{\min}^{2}}}\right)}\right)$-RDP and $2.1\alpha\varepsilon^{2}\left({\alpha+\frac{8}{\sigma_{\min}^{6}}}\right)$-RDP when coupled with PCA for rank reduction and standard scaling respectively.
@@
We apply Theorem 2 with $c_{1}=c_{2}=2$,
@@
By substituting the expression of RDP and SRDP parameter from Table 1 for Gaussian mechanism, the $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for deduplication, quantization, PCA, standard scaling and mean imputation from Proposition 3, 4, 5 and 6, and the Lipschitz parameter and global sensitivity of the output function $f$ into Equation 45, we complete the proof of the overall privacy guarantees for Gaussian mechanism.
@@
We apply Theorem 2 with $c_{1}=c_{2}=1$,
@@
Similarly, for exponential mechanism, we substitute the expression of RDP and SRDP parameter from Table 1 for exponential mechanism, the $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for deduplication, quantization, PCA, standard scaling and mean imputation from Proposition 3, 4, 5 and 6, and the Lipschitzness and the global sensitivity of the score function $Q$, into Equation 45. This completes the proof of the overall privacy guarantees for exponential mechanism.
@@
Following a similar argument as in the proof Gaussian mechanism (Theorem 17 (i)), we set the parameters in Theorem 2 $c_{1}=c_{2}=2$ and obtain Equation 45. Then, we substitute in Equation 45 the following parameters: a)RDP and SRDP parameters for DP-SGD with subsampling from Table 1, b)$\gamma=1$, c) Lipschitzness and smoothness parameter of the loss function $\ell$, and d) $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for PCA and standard scaling from Proposition 5 and 6. This completes the proof. ∎
For DP-SGD with iteration, the pre-processing mechanisms achieve a tighter bound while satisfying $\Delta_{\infty}\ll n$ by Definition 8. Therefore, this method will not provide tighter bound for PCA and standard scaling. However, it improves the privacy analysis for imputation, deduplication and quantization, as proved below.
@@
Following a similar argument as in the proof of Gaussian mechanism (Theorem 17 (i)), we set the parameters in Theorem 2 $c_{1}=c_{2}=2$. We substitute in Equation 45 the following parameters: a)RDP and SRDP parameters for DP-SGD with iteration from Table 1, b)$\kappa_{\tau}=\Delta_{\infty}=p$, c) Lipschitzness and smoothness parameter of the loss function $\ell$, and d) $L_{2}$ and $L_{\infty}$ sensitivity ($\Delta_{2}$ and $\Delta_{\infty}$) for quantization, deduplication and imputation from from Proposition 3 and Corollary 4. This completes the proof. ∎